OpenStudy (izzy18970):

I really need help on this. I don't get it!!!! Katie knocked down a quarter of the amount of pins Jacob knocked down in the bowling game. Fill in the box with the expression that represents the amount of pins Katie knocked down in the bowling game. Let k represent the amount of pins Katie knocked down and j represent the amount of pins Jacob knocked down. ___________ = k If Katie knocked down 84 pins, how many pins did Jacob knock down?

OpenStudy (anonymous):

1/4 x Jabob's pins = Katie's pins (1/4)j = k ---------

OpenStudy (anonymous):

That fills in the blank. Now, plug in 84 for k and solve for j to get the answer to the last question.

OpenStudy (izzy18970):

so j times 4 = k

OpenStudy (anonymous):

no. j times 1/4 = k

OpenStudy (izzy18970):

84 / 1/4 = 21

OpenStudy (izzy18970):

is this correct??